Make an equation of a line that passes through (2,-3) and is perpendicular to y=x-2
Found 2 solutions by fractalier, macston:
Answer by fractalier(6550)

If the line you want is perpendicular to y = x - 2, its slope must be -1.
Now plug into
y = mx + b and get
-3 = (-1)(2) + b
-3 = -2 + b
b = -1
and your line must be
y = -x - 1

Answer by macston(5194)

slope-intercept form:
y=mx+b
y=(1)x+(-2)
The slope is 1
The slope of the perpendicular line is the
negative inverse. Negative inverse of 1 is -1.
.
The equation will have the form:
y=(-1)x+b
Put in values of given point and solve for b:
-3=(-1)(2)+b
-3=-2+b
-1=b
Use this value of b in the equation:
y=(-1)x+(-1)
y=-x-1
.
ANSWER: Equation of the perpendicular line through (2,-3):
y=-x-1
.